Jack mowes the lawn in 4hrs and mary mowes the lawn in 7 hours. How long would it take them to both mow the yard?
Answer by checkley77(12844)

1/4+1/7=1/X
(7+4)/4*7=1/X
11/28=1/X CROSS MULTIPLY
11X=28
X=28/11
X=2.5454 HOURS WORKING TOGETHER.